OpenStudy (anonymous):

One number is 4 more than another, and their sum is 60. What is the smaller number? If x = the larger number and y = the smaller number, then which of the following systems could be used to solve the problem? a)x + y = 60 and -x + y = 4 b)x + y = 60 and x - y = 4 c)x + y = 60 and x + y = 4

OpenStudy (radar):

y=smaller number, given x=y+4, given y+y+4=60 (substituting y+4 for x), given 2y+4=60, 2y=56, y=28 then x=28+4=32 looking at the answers which are true? b